In the school's chemistry lab, a certain acid is stored in two different dilutons. One bottle contains a weaker solution that is only 25% acid, but another has a stronger solution of 65% acid. students must pour some from each container to make 60ml of a solution that is 50% acid. HOw much from each container should be used?
I know i'm going to use the two variables s-stronger and w-weaker. just can't figure out what the two equations going to be.

Answer by stanbon(57387)

In the school's chemistry lab, a certain acid is stored in two different dilutons. One bottle contains a weaker solution that is only 25% acid, but another has a stronger solution of 65% acid. students must pour some from each container to make 60ml of a solution that is 50% acid. HOw much from each container should be used?
-----
Equations:
Quantity: s + w = 60 ml
Acid::::0.65s+0.25w = 0.50*60
--------------------------------------
Multiply thru the Quantity equation by 65
Multiply thru the Acid equation by 100
--------------------------------------------
65s + 65w = 65*60
65s + 25w = 50*60
---------------------
subtract and solve for "w":
40w = 15*60
w = (3/2)15
w = 22.5 ml (amt. of weaker solution needed)
------
Solve for "s":
s + w = 60
s = 60-22.5 = 37.5 ml
------------------------------
Cheers,
Stan H.
===========================